OpenStudy (anonymous):

is (1,5) a solution to the equation y= x-2? Yes No

ganeshie8 (ganeshie8):

y = x - 2 plugin x = 1, do you get 5 ?

OpenStudy (anonymous):

5=1-2 . so no lol